OpenStudy (anonymous):

you are a judge at a air racing world championship. Three planes are in the race. Who is the winner??? The winner is determined by endurance and speed not by the first home.

OpenStudy (anonymous):

here are the details of the planes

OpenStudy (anonymous):

@Mertsj can you help???

OpenStudy (anonymous):

@dan815 could you help me?

OpenStudy (jack1):

can u upload as a pic / screenshot rather than a .doc dude?

OpenStudy (anonymous):

thanks @Jack1

OpenStudy (jack1):

in aviation endurance is literally : Max flight time for a given amount of fuel so i dont know how exactly to answer your question... because i'm not sure exactly what they're asking... hazarding a guess tho, and emphasis that this is only a GUESS ;P if we judge them on the aircraft's max speed, diligent should take 10 mins to complete the race, high flier should take 2.58 minutes to complete the race, and res should take 14 minutes to complete the race... now D took 20 HF took 35 and R took 28 so as a ratio: 20/10 = 2x expected time for Dil 35/2.578 = 13.58x expected time for HF 28/14 = 2x expected time for Res so I'd judge it to be a tie between Diligent and Resplendent... but that's just me assuming it's judged kinda like golf handicaps are...? lemme know how u go dude?

OpenStudy (anonymous):

thanks @Jack1 This has given me something to think about. Thanks once again for giving it a go!!